OpenStudy (anonymous):

What is -2x+2y=10 how do I solve this problem need HELP!

OpenStudy (anonymous):

Solving for what? x or y

OpenStudy (anonymous):

Solving the slope

OpenStudy (anonymous):

intercept

OpenStudy (anonymous):

rearrange in y = mx + c form So first isolate y on one side so we get 2y = 10 + 2x Divide both sides by 2 y = 5 + x y = x + 5

OpenStudy (anonymous):

So what's the intercept now? You should know that in y = mx + c, m = gradient and c = y intercept

OpenStudy (anonymous):

y=mx+b